# The asymptotic Plancherel formula and Lusztig’s asymptotic algebra for $\tilde{\mathsf{A}}_{n}$

### v1

@@ 2.1 The $\lambda$-weights and the fundamental $\lambda$-alcove
By [10, Lemma 2.3], for each $\gamma\in P$ there exists a unique element $\gamma^{(\lambda)}\in\mathcal{A}_{\lambda}\cap P$ with $\gamma-\gamma^{(\lambda)}\in Q_{\lambda}$ (we call $\gamma^{(\lambda)}$ the projection of $\gamma$ onto the fundamental $\lambda$-alcove). The set
@@
is closely related to the set $P/Q_{\lambda}$ of $\lambda$-weights (see Proposition 2.6 below). By [10, Lemma 2.5] an element $\gamma=\sum_{i=1}^{n}a_{i}\omega_{i}\in P$ lies in $P^{(\lambda)}$ if and only if whenever $j\in J_{\lambda}$ we have $a_{j}\in\{0,1\}$, and for each connected component $K$ of $J_{\lambda}$ there is at most one $k\in K$ with $a_{k}=1$ (the connected components of $J_{\lambda}$ are the sets $\{\lambda(i-1)+1,\ldots,\lambda(i)-1\}$ for $1\leq i\leq r(\lambda)$).
@@
If $\Phi$ is of type $\mathsf{A}_{2}$ and $\lambda=(2,1)$ then $\mathcal{A}_{\lambda}$ (shaded green) and $P^{(\lambda)}$ (represented as bullets) are as follows.
@@
The elements $\tau_{\gamma}$ are called pseudo-translations, and by [10, Lemma 2.12] we have
By [10, Lemma 2.8] we have $\Phi(\mathsf{y}_{\gamma})=\{\alpha\in\Phi_{\lambda}^{+}\mid\langle\gamma,\alpha\rangle=1\}$. Note that $\tau_{\gamma}$ and $\mathsf{y}_{\gamma}$ depend on the partition $\lambda$, however this dependence is suppressed without risk of confusion.
Proposition 2.5 ([10, Corollary 2.13]).
@@
Thus $P^{(\lambda)}$ inherits the structure of an abelian group with operation $\gamma\oplus\gamma^{\prime}=(\gamma+\gamma^{\prime})^{(\lambda)}$, and $P^{(\lambda)}\cong P/Q_{\lambda}$ via the map $\gamma\mapsto\gamma+Q_{\lambda}$. The following description of the inverse isomorphism $P/Q_{\lambda}\to P^{(\lambda)}$ is very useful (for example, combined with (2.1) it gives an explicit expression for the projection map $P\to P^{(\lambda)}$). Recall that $\lambda[i,j]$ denotes the element in row $i$ and column $j$ of $\mathbf{t}_{r}(\lambda)$.
@@ 2.2 The group $G_{\lambda}$
be the subgroup of $W$ stabilising the fundamental $\lambda$-alcove. By [10, Theorem 2.18] we have
@@
Moreover $g\in G_{\lambda}$ maps the simple roots of $\Phi_{\lambda}$ to the simple roots of $\Phi_{\lambda}$, and hence induces a permutation of $J_{\lambda}$. This permutation maps connected components of $J_{\lambda}$ to connected components of $J_{\lambda}$. Furthermore, by [10, Lemma 2.19] for $\gamma\in P^{(\lambda)}$ and $g\in G_{\lambda}$ we have
@@
The subgroup of $\widetilde{W}$ stabilising $\mathcal{A}_{\lambda}$ is (see [10, Corollary 2.20])
@@
We have
$G_{\lambda}=\{g\in\mathsf{u}_{\lambda}W_{\lambda^{\prime}}\mathsf{u}_{\lambda}^{-1}\mid g\Phi_{\lambda}=\Phi_{\lambda}\}.$
In particular, $\mathsf{u}_{\lambda}^{-1}G_{\lambda}\mathsf{u}_{\lambda}\leq W_{\lambda^{\prime}}$.
@@ 2.3 Dominant $\lambda$-weights and the $\lambda$-dominance order
Let $\gamma=\sum_{i=1}^{n+1}a_{i}e_{i}\in Q$, with the expression chosen so that $a_{1}+\cdots+a_{n+1}=0$. Then $\gamma\in Q_{\lambda}$ if and only if
$\sum_{i=\lambda(k-1)+1}^{\lambda(k)}a_{i}=0\quad\text{ for each $1\leq k\leq r(\lambda)$}.$
@@
Let $\lambda=(6,6,4,4,4,2,1,1)$ as in Example 2.9. Then $Q^{\lambda}_{+}$ consists of the elements $a_{1}\tilde{e}_{1}+\cdots+a_{8}\tilde{e}_{8}$ with $a_{i}\in\mathbb{Z}$ satisfying
$a_{1}+a_{2}=a_{3}+a_{4}+a_{5}=a_{6}=a_{7}+a_{8}=0,\quad a_{1}\geq 0,\quad a_{3}\geq 0,\quad a_{3}+a_{4}\geq 0,\quad a_{7}\geq 0.$
@@ 2.5 The ring $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$ and $G_{\lambda}$-Schur functions
A more explicit description of the ring $\mathsf{R}[\zeta_{\lambda}]$ is as follows. Since $\zeta_{\lambda}^{\alpha}=1$ for all $\alpha\in\Phi_{\lambda}$ we have $\zeta_{j}+\mathcal{I}_{\lambda}=\zeta_{j^{\prime}}+\mathcal{I}_{\lambda}$ whenever $j$ and $j^{\prime}$ lie in the same row of $\mathbf{t}_{r}(\lambda)$. Define $z_{1},\ldots,z_{r(\lambda)}\in\mathsf{R}[\zeta_{\lambda}]$ by $z_{i}=\zeta_{j}+\mathcal{I}_{\lambda}$ for any $j$ in the $i$th row of $\mathbf{t}_{r}(\lambda)$ (that is, $z_{i}=\zeta_{\lambda}^{e_{j}}$). Then the relation $e_{1}+\cdots+e_{n+1}=0$ gives
$z_{1}^{\lambda_{1}}z_{2}^{\lambda_{2}}\cdots z_{r(\lambda)}^{\lambda_{r(\lambda)}}=1,$
and $\mathsf{R}[\zeta_{\lambda}]$ may be regarded as the Laurent polynomial ring in the indeterminates $z_{1}^{\pm 1},\ldots,z_{r(\lambda)}^{\pm 1}$ subject to the above equation.
@@
Let
$\mathsf{R}[\zeta_{\lambda}]^{G_{\lambda}}=\{p(\zeta_{\lambda})\in\mathsf{R}[\zeta_{\lambda}]\mid g\cdot p(\zeta_{\lambda})=p(\zeta_{\lambda})\text{ for all $g\in G_{\lambda}$}\}$
and similarly $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}=\{p(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]\mid g\cdot p(\zeta_{\lambda})=p(\zeta_{\lambda})\text{ for all $g\in G_{\lambda}$}\}$.
@@
This is classical, see for example, [23, (2.14)]. ∎
@@
Recall that $z_{1}^{2}z_{2}^{2}=1$ (however note that $z_{1}z_{2}\neq 1$). Thus, in particular, $\mathfrak{s}_{\tilde{e}_{1}+\tilde{e}_{2}}(\zeta_{\lambda})=z_{1}z_{2}$. Note that $\mathfrak{s}_{\tilde{e}_{1}+\tilde{e}_{2}}(\zeta_{\lambda})^{2}=z_{1}^{2}z_{2}^{2}=1$.
@@ 2.6 Representations of $\widetilde{H}$ induced from Levi subalgebras
$\psi_{\lambda}(T_{j})$ $=\mathsf{v}$ for $j\in J_{\lambda}$
$\psi_{\lambda}(X^{\gamma})$ $=\mathsf{v}^{\langle\gamma,2\rho_{\lambda}\rangle}\zeta_{\lambda}^{\gamma}$ for $\gamma\in P$
uniquely extends to a multiplicative character $\psi_{\lambda}:\mathcal{L}_{\lambda}\to\mathsf{R}[\zeta_{\lambda}]$ (see [10, §4.2]).
@@
By [10, Proposition 4.20] $M_{\lambda}$ is a free $\mathsf{R}[\zeta_{\lambda}]$-module with basis
@@
By [10, Theorem 4.13] we have the following combinatorial formula.
Theorem 2.28 ([10, Theorem 4.13]).
@@ 4. $(4)$
(1), (2) and (3) are direct calculations using the Bernstein relation, and (4) follows by induction on $\ell(v)$. ∎
@@ 3.3 $G_{\lambda}$ symmetry of $f_{\lambda}(h)$
Each element of the Hecke algebra can be written as a linear combination of the elements $U_{w}$, $w\in W$, with “coefficients” being rational functions (in the variables $X^{\gamma}$) whose denominators do not vanish on applying $\psi_{\lambda}$. Therefore, by linearity of $f_{\lambda}$, it is sufficient to prove that $f_{\lambda}(p(X)U_{v})$ is $G_{\lambda}$-symmetric, where $v\in W$ and where $p(X)$ is a rational function of the form described above. By definition we have
$f_{\lambda}(p(X)U_{v})=\chi_{\lambda}(p(X)U_{v}C_{\mathsf{w}_{\lambda^{\prime}}}),$
and we shall compute this character below.
@@ 3.4 $\lambda$-relative Satake isomorphism
With this normalisation we have $\mathbf{1}_{\lambda^{\prime}}^{2}=\mathbf{1}_{\lambda^{\prime}}$, and the algebra
$\pi_{\lambda}(\mathbf{1}_{\lambda^{\prime}}\widetilde{H}_{\mathsf{R}^{\prime}}\mathbf{1}_{\lambda^{\prime}})$
is unital (with identity $\mathbf{1}_{\lambda^{\prime}}$) and commutative (by Corollary 3.4).
@@ 4.1 The killing property
In the basis of intertwiners, the matrices $\pi_{\lambda}(U_{w})$ (for $w\in W$) have distinct support (the places of the nonzero entries; because the $u$th row of $\pi_{\lambda}(U_{w})$ is either 00, or has an entry only in the $uw$ position in the case $uw\in{{}^{\lambda}}W$, see Proposition 2.30). Since $\pi_{\lambda}(c_{w}^{\mu^{\prime}}(X))$ is diagonal (in the intertwiner basis) it follows that $\pi_{\lambda}(C_{\mathsf{w}_{\mu^{\prime}}})=0$ if and only if
$\pi_{\lambda}(c_{w}^{\mu^{\prime}}(X)U_{w})=0\quad\text{for all $w\in W_{\mu^{\prime}}$}.$
This, in turn, is equivalent to the statement
@@ 4.2 Boundedness
Let $N=\max\{\deg[\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})]_{u,v}\mid w\in\widetilde{W},\,u,v\in{{}^{\lambda}}W\}$ which is well defined by [10, Theorem 5.13]. Let $w$ be such that $\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})$ attains degree $N$, and suppose that this degree is attained on the $u$th row of $\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})$. Let $v_{1},\ldots,v_{k}\in{{}^{\lambda}}W$ be the columns with $\deg[\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})]_{u,v_{j}}=N$ for $1\leq j\leq k$. Write
@@
where the omitted terms in the sum are of strictly lower degree. We claim that the coefficient of $\mathsf{q}^{2N}$ cannot vanish. To see this, note that if $a(\zeta_{\lambda})\in\mathsf{R}[\zeta_{\lambda}]$ with $a(\zeta_{\lambda})=\sum_{\gamma\in P^{(\lambda)}}a_{\gamma}\zeta_{\lambda}^{\gamma}$ with $a_{\gamma}\in\mathbb{Z}$ then
$a(\zeta_{\lambda})a(\zeta_{\lambda}^{-1})=\sum_{\gamma_{1},\gamma_{2}}a_{\gamma_{1}}a_{\gamma_{2}}\zeta_{\lambda}^{\gamma_{1}-\gamma_{2}}=\sum_{\gamma}a_{\gamma}^{2}+\text{terms involving $\zeta_{\lambda}$}.$
In particular, the constant term is strictly positive. Since $[a_{1}(\zeta_{\lambda})a_{1}(\zeta_{\lambda}^{-1})+\cdots+a_{k}(\zeta_{\lambda})a_{k}(\zeta_{\lambda}^{-1})]$ is a sum of terms of this form, it cannot vanish. In summary, $\pi_{\lambda}(C_{w};\mathsf{B}_{\lambda})\pi_{\lambda}(C_{w^{-1}};\mathsf{B}_{\lambda})$ attains degree $2N$ in the $(u,u)$-entry.
@@
The following corollary verifies [10, Conjecture 5.16] for type $\tilde{\mathsf{A}}_{n}$.
@@
4.3 Leading matrices

### v2

@@ 2.1 The $\lambda$-weights and the fundamental $\lambda$-alcove
By [13, Lemma 2.3], for each $\gamma\in P$ there exists a unique element $\gamma^{(\lambda)}\in\mathcal{A}_{\lambda}\cap P$ with $\gamma-\gamma^{(\lambda)}\in Q_{\lambda}$ (we call $\gamma^{(\lambda)}$ the projection of $\gamma$ onto the fundamental $\lambda$-alcove). The set
@@
is closely related to the set $P/Q_{\lambda}$ of $\lambda$-weights (see Proposition 2.6 below). By [13, Lemma 2.5] an element $\gamma=\sum_{i=1}^{n}a_{i}\omega_{i}\in P$ lies in $P^{(\lambda)}$ if and only if whenever $j\in J_{\lambda}$ we have $a_{j}\in\{0,1\}$, and for each connected component $K$ of $J_{\lambda}$ there is at most one $k\in K$ with $a_{k}=1$ (the connected components of $J_{\lambda}$ are the sets $\{\lambda(i-1)+1,\ldots,\lambda(i)-1\}$ for $1\leq i\leq r(\lambda)$).
The elements $\tau_{\gamma}$ are called pseudo-translations, and by [13, Lemma 2.12] we have
By [13, Lemma 2.8] we have $\Phi(\mathsf{y}_{\gamma})=\{\alpha\in\Phi_{\lambda}^{+}\mid\langle\gamma,\alpha\rangle=1\}$. Note that $\tau_{\gamma}$ and $\mathsf{y}_{\gamma}$ depend on the partition $\lambda$, however this dependence is suppressed without risk of confusion.
@@
If $\Phi$ is of type $\mathsf{A}_{2}$ and $\lambda=(2,1)$ then $\mathcal{A}_{\lambda}$ (shaded green) and $P^{(\lambda)}$ (represented as bullets) are as illustrated in Figure 1. If $\gamma\in P^{(\lambda)}$ lies on the hyperplane $H_{\alpha_{1},1}$ then $\mathsf{y}_{\gamma}=s_{1}$.
@@
Proposition 2.5 ([13, Corollary 2.13]).
@@
Thus $P^{(\lambda)}$ inherits the structure of an abelian group with operation $\gamma\oplus\gamma^{\prime}=(\gamma+\gamma^{\prime})^{(\lambda)}$, and $P^{(\lambda)}\cong P/Q_{\lambda}$ via the map $\gamma\mapsto\gamma+Q_{\lambda}$. The following description of the inverse isomorphism $P/Q_{\lambda}\to P^{(\lambda)}$ is useful (combined with (2.1) it gives an explicit expression for the projection map $P\to P^{(\lambda)}$). Recall that $\lambda[i,j]$ denotes the element in row $i$ and column $j$ of $\mathbf{t}_{r}(\lambda)$.
@@ 2.2 The group $G_{\lambda}$
be the subgroup of $W$ stabilising the fundamental $\lambda$-alcove. By [13, Theorem 2.18] we have
@@
Moreover $g\in G_{\lambda}$ maps the simple roots of $\Phi_{\lambda}$ to the simple roots of $\Phi_{\lambda}$, and hence induces a permutation of $J_{\lambda}$. This permutation maps connected components of $J_{\lambda}$ to connected components of $J_{\lambda}$. Furthermore, by [13, Lemma 2.19] for $\gamma\in P^{(\lambda)}$ and $g\in G_{\lambda}$ we have
@@
The subgroup of $\widetilde{W}$ stabilising $\mathcal{A}_{\lambda}$ is (see [13, Corollary 2.20])
@@
We have $G_{\lambda}=\{g\in\mathsf{u}_{\lambda}W_{\lambda^{\prime}}\mathsf{u}_{\lambda}^{-1}\mid g\Phi_{\lambda}=\Phi_{\lambda}\}.$ In particular, $\mathsf{u}_{\lambda}^{-1}G_{\lambda}\mathsf{u}_{\lambda}\leq W_{\lambda^{\prime}}$.
@@ 2.3 Dominant $\lambda$-weights and the $\lambda$-dominance order
Let $\gamma=\sum_{i=1}^{n+1}a_{i}e_{i}\in Q$, with the expression chosen so that $a_{1}+\cdots+a_{n+1}=0$. Then $\gamma\in Q_{\lambda}$ if and only if $\sum_{i=\lambda(k-1)+1}^{\lambda(k)}a_{i}=0$ for each $1\leq k\leq r(\lambda)$.
@@
Let $\lambda=(6,6,4,4,4,2,1,1)$ as in Example 2.9. Then $Q^{\lambda}_{+}$ consists of the elements $a_{1}\tilde{e}_{1}+\cdots+a_{8}\tilde{e}_{8}$ with $a_{i}\in\mathbb{Z}$ satisfying $a_{1}+a_{2}=a_{3}+a_{4}+a_{5}=a_{6}=a_{7}+a_{8}=0$, $a_{1}\geq 0$, $a_{3}\geq 0$, $a_{3}+a_{4}\geq 0$, and $a_{7}\geq 0$.
@@ 2.5 The ring $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$ and $G_{\lambda}$-Schur functions
A more explicit description of the ring $\mathsf{R}[\zeta_{\lambda}]$ is as follows. Since $\zeta_{\lambda}^{\alpha}=1$ for all $\alpha\in\Phi_{\lambda}$ we have $\zeta_{j}+\mathcal{I}_{\lambda}=\zeta_{j^{\prime}}+\mathcal{I}_{\lambda}$ whenever $j$ and $j^{\prime}$ lie in the same row of $\mathbf{t}_{r}(\lambda)$. Define $z_{1},\ldots,z_{r(\lambda)}\in\mathsf{R}[\zeta_{\lambda}]$ by $z_{i}=\zeta_{j}+\mathcal{I}_{\lambda}$ for any $j$ in the $i$th row of $\mathbf{t}_{r}(\lambda)$ (that is, $z_{i}=\zeta_{\lambda}^{e_{j}}$). Then the relation $e_{1}+\cdots+e_{n+1}=0$ gives $z_{1}^{\lambda_{1}}z_{2}^{\lambda_{2}}\cdots z_{r(\lambda)}^{\lambda_{r(\lambda)}}=1,$ and $\mathsf{R}[\zeta_{\lambda}]$ may be regarded as the Laurent polynomial ring in the indeterminates $z_{1}^{\pm 1},\ldots,z_{r(\lambda)}^{\pm 1}$ subject to the above equation.
@@
Let $\mathsf{R}[\zeta_{\lambda}]^{G_{\lambda}}=\{p(\zeta_{\lambda})\in\mathsf{R}[\zeta_{\lambda}]\mid g\cdot p(\zeta_{\lambda})=p(\zeta_{\lambda})\text{ for all $g\in G_{\lambda}$}\}$ and similarly $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}=\{p(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]\mid g\cdot p(\zeta_{\lambda})=p(\zeta_{\lambda})\text{ for all $g\in G_{\lambda}$}\}$.
@@
This is classical, see for example, [26, (2.14)]. ∎
@@
Recall that $z_{1}^{2}z_{2}^{2}=1$. In particular $\mathfrak{s}_{\tilde{e}_{1}+\tilde{e}_{2}}(\zeta_{\lambda})=z_{1}z_{2}$, and $\mathfrak{s}_{\tilde{e}_{1}+\tilde{e}_{2}}(\zeta_{\lambda})^{2}=z_{1}^{2}z_{2}^{2}=1$.
@@ 2.6 Representations of $\widetilde{H}$ induced from Levi subalgebras
$\psi_{\lambda}(T_{j})=\mathsf{v}\quad\text{and}\quad\psi_{\lambda}(X^{\gamma})=\mathsf{v}^{\langle\gamma,2\rho_{\lambda}\rangle}\zeta_{\lambda}^{\gamma}$
for $j\in J_{\lambda}$ and $\gamma\in P$ extends to a multiplicative character $\psi_{\lambda}:\mathcal{L}_{\lambda}\to\mathsf{R}[\zeta_{\lambda}]$ (see [13, §4.2]).
@@
By [13, Proposition 4.20] $M_{\lambda}$ is a free $\mathsf{R}[\zeta_{\lambda}]$-module with basis
@@
By [13, Theorem 4.13] we have the following combinatorial formula.
Theorem 2.28 ([13, Theorem 4.13]).
@@ 4. $(4)$
(1), (2), (3) follow from the Bernstein relation, and (4) follows by induction on $\ell(v)$. ∎
@@ 3.3 $G_{\lambda}$ symmetry of $f_{\lambda}(h)$
Each element of the Hecke algebra can be written as a linear combination of the elements $U_{w}$, $w\in W$, with “coefficients” being rational functions (in the variables $X^{\gamma}$) whose denominators do not vanish on applying $\psi_{\lambda}$. Therefore, by linearity of $f_{\lambda}$, it is sufficient to prove that $f_{\lambda}(p(X)U_{v})$ is $G_{\lambda}$-symmetric, where $v\in W$ and where $p(X)$ is a rational function of the form described above. By definition we have $f_{\lambda}(p(X)U_{v})=\chi_{\lambda}(p(X)U_{v}C_{\mathsf{w}_{\lambda^{\prime}}}),$ and we shall compute this character below.
@@ 3.4 $\lambda$-relative Satake isomorphism
With this normalisation we have $\mathbf{1}_{\lambda^{\prime}}^{2}=\mathbf{1}_{\lambda^{\prime}}$, and the algebra $\pi_{\lambda}(\mathbf{1}_{\lambda^{\prime}}\widetilde{H}_{\mathsf{R}^{\prime}}\mathbf{1}_{\lambda^{\prime}})$ is unital (with identity $\mathbf{1}_{\lambda^{\prime}}$) and commutative (by Corollary 3.4).
@@ 4.1 The killing property
In the basis of intertwiners, the matrices $\pi_{\lambda}(U_{w})$ (for $w\in W$) have distinct support (the places of the nonzero entries; because the $u$th row of $\pi_{\lambda}(U_{w})$ is either 00, or has an entry only in the $uw$ position in the case $uw\in{{}^{\lambda}}W$, see Proposition 2.30). Since $\pi_{\lambda}(c_{w}^{\mu^{\prime}}(X))$ is diagonal (in the intertwiner basis) it follows that $\pi_{\lambda}(C_{\mathsf{w}_{\mu^{\prime}}})=0$ if and only if $\pi_{\lambda}(c_{w}^{\mu^{\prime}}(X)U_{w})=0$ for all $w\in W_{\mu^{\prime}}$. This, in turn, is equivalent to the statement
@@ 4.2 Boundedness
Let $N=\max\{\deg[\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})]_{u,v}\mid w\in\widetilde{W},\,u,v\in{{}^{\lambda}}W\}$ which is well defined by [13, Theorem 5.13]. Let $w$ be such that $\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})$ attains degree $N$, and suppose that this degree is attained on the $u$th row of $\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})$. Let $v_{1},\ldots,v_{k}\in{{}^{\lambda}}W$ be the columns with $\deg[\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})]_{u,v_{j}}=N$ for $1\leq j\leq k$. Write
@@
where the omitted terms in the sum are of strictly lower degree. We claim that the coefficient of $\mathsf{q}^{2N}$ cannot vanish. To see this, note that if $a(\zeta_{\lambda})\in\mathsf{R}[\zeta_{\lambda}]$ with $a(\zeta_{\lambda})=\sum_{\gamma\in P^{(\lambda)}}a_{\gamma}\zeta_{\lambda}^{\gamma}$ with $a_{\gamma}\in\mathbb{Z}$ then $a(\zeta_{\lambda})a(\zeta_{\lambda}^{-1})=\sum_{\gamma_{1},\gamma_{2}}a_{\gamma_{1}}a_{\gamma_{2}}\zeta_{\lambda}^{\gamma_{1}-\gamma_{2}}=\sum_{\gamma}a_{\gamma}^{2}+\text{terms involving $\zeta_{\lambda}$}.$ In particular, the constant term is strictly positive. Since $[a_{1}(\zeta_{\lambda})a_{1}(\zeta_{\lambda}^{-1})+\cdots+a_{k}(\zeta_{\lambda})a_{k}(\zeta_{\lambda}^{-1})]$ is a sum of terms of this form, it cannot vanish. In summary, $\pi_{\lambda}(C_{w};\mathsf{B}_{\lambda})\pi_{\lambda}(C_{w^{-1}};\mathsf{B}_{\lambda})$ attains degree $2N$ in the $(u,u)$-entry.
@@
The following corollary verifies [13, Conjecture 5.16] for type $\tilde{\mathsf{A}}_{n}$.
@@
4.3 Leading matrices and the characters $\chi_{\lambda}^{\infty}$ of $\mathcal{J}$